# Gromov-Hausdorff limits of smocked spaces

### v2

@@ Abstract
Smocked spaces are a class of metric spaces which were introduced to generalise pulled thread spaces. We investigate convergence of these spaces, showing that the smocked space obtained from the Hausdorff limit of a sequence of smocking sets is equivalent to the Gromov-Hausdorff limit of the corresponding smocked spaces. We prove that it is sufficient to have uniform bounds on the smocking constants in order to have a sequence of smocked spaces with a Gromov-Hausdorff converging subsequence. We also show that given a normed space, there exists a smocked space whose unique tangent cone at infinity is that space. This answers three open questions of Sormani et al. We finish with some remarks on possible applications to metric measure spaces.
@@ 1 Introduction
The article is structured as follows. In Section 2, we state basic definitions and results. In Section 3, we prove that if one defines a sequence of smocked spaces using a sequence of smocking sets which has Hausdorff convergence to a limiting smocking set, then the corresponding smocked spaces converge in the pointed Gromov-Hausdorff sense. We also investigate sufficient conditions for a sequence of smocked spaces to have a subsequence which Gromov-Hausdorff converges, finding that it is sufficient to have uniform bounds on the smocking constants. Furthermore, we show that given a finite-dimensional normed space, there exists a smocked space with a unique tangent cone at infinity which is that space. Section 4 is somewhat more conjectural. We consider the possibility of extending to smocked spaces equipped with measures and also smocked sets which are defined for Wasserstein spaces of measures.
@@ 2.2 Smocked spaces
$\pi(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\in\mathbb{E}^{n}\>\backslash\>S\\ I_{j}&\mbox{if }x\in I_{j},\>j\in J\\ \end{array}\right.$
@@
The definition of a smocked space refers to a set of four constants, which are referred to in [2] as the smocking constants: these are the smocking depth $h$, the maximum smocking length $L_{\text{max}}$, the minimum smocking length $L_{\text{min}}$, and the smocking separation factor $\delta$, which was already defined in Definition ?.
@@ 2.3 Metric geometry
To any metric space $X$, one may also canonically associate another metric space called the Wasserstein space [14]. These spaces will be mentioned in Section 4.
Definition 2.10: The Wasserstein space $\mathcal{P}_{p}(X):=(P(X),W_{p})$ is a metric space associated to $X$, where $P(X)$ is the space of Borel probability measures on $X$ and $W_{p}$ is the Wasserstein distance of order $p$ between two measures $\mu$ and $\nu$, which is defined to be
@@
where $\pi$ can take values in the set of all transference plans between $\mu$ and $\nu$.
@@ 3.1 Relation between Hausdorff and Gromov-Hausdorff convergence
We start with a result which relates Hausdorff convergence of smocking sets to Gromov-Hausdorff convergence of the corresponding smocked spaces.
Proposition 3.1: If one defines a sequence of smocked spaces $X_{k}$ using smocking sets $S_{k}\subset\mathbb{E}^{N}$ such that the sets Hausdorff converge in the limit to a smocking set $S_{\infty}$, then the corresponding smocked spaces converge to $X_{\infty}$ in the pointed Gromov-Hausdorff sense.
Proof: Smocked spaces are complete and noncompact (see Theorem 6.1 of [2]), so we use pointed Gromov-Hausdorff convergence as defined in Section 2. We must show that for all $r>0$
$\lim_{k\rightarrow\infty}d_{GH}\big{(}\big{(}B_{r}(x_{0}^{(k)}),d_{X_{k}}\big{)},\big{(}B_{r}(x_{0}^{(\infty)}),d_{X_{\infty}}\big{)}\big{)}=0,$
where $X_{\infty}$ is the smocked space obtained from the Hausdorff limit $S_{\infty}$ and $x_{0}^{(k)}$ is a base point in $X_{k}$.
Consider the smocking maps $\pi_{k}:\mathbb{E}^{N}\rightarrow X_{k}$ and $\pi_{\infty}:\mathbb{E}^{N}\rightarrow X_{\infty}$ associated to $X_{k}$ and $X_{\infty}$, respectively. By shifting the smocking set $S_{k}$, we can assume that
$\pi_{k}(0)=x_{0}^{(k)}\>\>\>\>\>\>\text{for any}\>\pi_{k},$
with a similar statement holding for $S_{\infty}$. We then define a subset $U_{r}(x_{0}^{(k)})$ of $\mathbb{E}^{N}$ defined for any $r>0$:
$U_{r}(x_{0}^{(k)})=\pi_{k}^{-1}(B_{r}(x_{0}^{(k)}))\subset\mathbb{E}^{N}.$
As our correspondence for each $r>0$, we take
$C=\{(\pi_{k}(w),\pi_{\infty}(w)):w\in U_{r}(x_{0}^{(k)})\}\subset B_{r}(x_{0}^{(k)})\times B_{r}(x_{0}^{(\infty)}).$
This is a correspondence because both smocking maps are surjections. Next, note that
$|d_{X_{k}}(\pi_{k}(v),\pi_{k}(w))-d_{X_{\infty}}(\pi_{\infty}(v),\pi_{\infty}(w))|=|\overline{d}_{k}(v,w)-\overline{d}_{\infty}(v,w)|.$
The collection of sets $S_{\infty}$ is a Hausdorff limit, so it is the collection of sets of limits of all converging sequences $\{s_{k}\}$ in $\mathbb{E}^{N}$ such that $s_{k}\in S_{k}$ for all $k$. In particular, such sequences can only grow with increasing $k$ with at most exponential speed, so the distances measured by the respective smocking pseudometrics can only decay with at most exponential speed. This implies that we have for the difference of the distances
$|\overline{d}_{k}(v,w)-\overline{d}_{\infty}(v,w)|\leq Ae^{-f(k)},$
where $A$ is some sufficiently large constant and $f(k)$ is a monotonically increasing function of $k$. On the other hand, such sequences could instead decay with at most exponential speed, and the same conclusion holds after taking the magnitude of the difference. By Theorem 7.3 of [2] for almost distance preserving correspondences, this implies that
$d_{GH}\big{(}\big{(}B_{r}(x_{0}^{(k)}),d_{X_{k}}\big{)},\big{(}B_{r}(x_{0}^{(\infty)}),d_{X_{\infty}}\big{)}\big{)}\leq 2Ae^{-f(k)}.$
Taking the limit, we have
$\lim_{k\rightarrow\infty}d_{GH}\big{(}\big{(}B_{r}(x_{0}^{(k)}),d_{X_{k}}\big{)},\big{(}B_{r}(x_{0}^{(\infty)}),d_{X_{\infty}}\big{)}\big{)}=0.$
@@
Some explicit examples may help to illustrate this proposition:
$X_{k}=\{x\in\mathbb{E}^{N}\backslash\overline{B}_{1/k}(0)\}\cup\{\overline{B}_{1/k}(0)\}.$
@@
as required. In general, the Gromov-Hausdorff limit $X_{\infty}$ will not necessarily be $\mathbb{E}^{N}$ and there will be a possibly countable collection of disjoint compact connected sets in each smocking set in the sequence, so it was necessary to show that a formal argument with correspondences goes through in the general case.
It is possible to think of potential counterexamples which are something as follows: one imagines a smocking pattern given by alternating balls of radius $1/k$ with balls of radius $(2^{k}-2)/2^{k}$. In the case of $k=1$, this implies that $S_{1}$ is a pattern which alternates between unit $N$-balls and points. In the limit as $k$ goes to infinity, $S_{k}$ will become the opposite pattern, since the unit balls in $S_{1}$ will all shrink to points and the points will expand out to be unit balls. This will give Hausdorff convergence to a smocking set $S_{\infty}$ which can be smocked to obtain a well-defined smocked space $X_{\infty}$.
However, $(2^{k}-2)/2^{k}$ approaches $1$ faster than $1/k$ approaches zero as $k$ goes to infinity. It follows that if the balls in $S_{1}$ are sufficiently close to the points, at some stage in the sequence of smocking sets $S_{k}$ the growing ball will merge with the shrinking ball to form a new set which might develop such that the difference between the distance between two points in $X_{k}$ and the distance between the corresponding two points in $X_{k+1}$ at some stage momentarily gets larger, whereas one would expect the difference between distances between pairs of points to decrease with increasing $k$ if there is to be Gromov-Hausdorff convergence of the metric spaces as $k$ tends to infinity. This fails to be a counterexample because by hypothesis the family of subsets over $\mathbb{E}^{N}$ need to remain disjoint from each other as the sequence $\{S_{k}\}_{k=1}^{\infty}$ converges to its Hausdorff limit.

### v3

# Gromov–Hausdorff convergence and tangent cones of smocked spaces

@@ Abstract
Smocked spaces are a class of metric spaces which were introduced to generalize pulled thread spaces. We investigate convergence of these spaces, showing that if the underlying smocking sets converge in Hausdorff distance and satisfy local uniform bounds on the smocking constants, then the associated smocked spaces converge in the pointed Gromov-Hausdorff sense. We prove a corresponding precompactness result using a similar assumption on the smocking constants. We also show that every finite-dimensional normed vector space arises as the tangent cone at infinity of a suitably constructed smocked space. Finally, we extend the convergence theory to the setting of smocked metric measure spaces, proving stability under pointed measured Gromov-Hausdorff convergence. These results establish a basic geometric framework for smocked spaces and demonstrate that they exhibit controlled limit behavior.
Keywords: Smocked spaces, Gromov–Hausdorff convergence, metric geometry, Hausdorff convergence, tangent cones, metric measure spaces.
53C23, 54E35; Secondary 28A33, 51F30.
@@ 1 Introduction
The article is structured as follows. In Section 2, we state basic definitions and results. In Section 3, we give a counterexample which shows that Hausdorff convergence of a sequence of smocking sets $S_{k}$ to a limit set $S_{\infty}$ does not necessarily imply that the corresponding smocked spaces $X_{k}$ converge to $X_{\infty}$ in the pointed Gromov-Hausdorff sense, answering an open question of Sormani et al. in the negative. We show that the smocked spaces do converge if one assumes local Hausdorff convergence on balls and local control of the smocking constants used to define a smocking space. We show that a similar story also holds for subsequential Gromov-Hausdorff convergence: global bounds on the smocking constants are sufficient for precompactness, but one must assume stronger local uniform bounds in order to have necessity. We end the section by showing that for any given finite-dimensional normed space, there is always a smocked space whose unique tangent cone at infinity is that space. In Section 4, we show that these convergence results extend naturally to smocked spaces equipped with measures. We finish with conclusions and future directions in Section 5.
@@ 2.2 Smocked spaces
[…]
@@
The definition of a smocked space refers to a set of four constants, which are referred to in [2] as the smocking constants: these are the smocking depth $h$, the maximum smocking length $L_{\text{max}}$, the minimum smocking length $L_{\text{min}}$, and the smocking separation factor $\delta$, which was already defined in Definition 2.6.
@@ 2.3 Metric geometry
We will also use two additional elementary lemmas which we prove here (these are used implicitly in the proof of Proposition 3.3).
Lemma 2.10 Let $\{S_{\alpha}\}$ be a collection of pairwise disjoint compact subsets (stitches) of $\overline{B_{R+1}(0)}\subset\mathbb{E}^{N}$ such that each $S_{\alpha}$ has diameter $\leq L$ and distinct stitches are separated by distance at least $\delta_{0}>0$. Then any rectifiable curve $\tilde{\gamma}\subset\overline{B_{R+1}(0)}$ of Euclidean length $\ell(\tilde{\gamma})\leq L_{0}$ meets at most
$M\;:=\;1+\Big\lfloor\frac{L_{0}}{\delta_{0}}\Big\rfloor$
distinct stitches. In particular, for $L_{0}=2R+2$ the number of stitches intersected is bounded by a constant $M(R,L,\delta_{0})$ depending only on $R,L,\delta_{0}$.
Proof: Order the distinct stitches met by $\tilde{\gamma}$ in the sequence encountered along the curve $S_{\alpha_{1}},S_{\alpha_{2}},\dots,S_{\alpha_{m}}$. For each consecutive pair $S_{\alpha_{j}},S_{\alpha_{j+1}}$ choose points $p_{j}\in S_{\alpha_{j}}$ and $q_{j}\in S_{\alpha_{j+1}}$ which lie along the curve in that order. By the separation hypothesis, we assume that $|p_{j}-q_{j}|_{\mathbb{E}^{N}}\geq\delta_{0}$. Summing over the m1m-1 gaps, we obtain
$\ell(\tilde{\gamma})\ \geq\ \sum_{j=1}^{m-1}|p_{j}-q_{j}|\ \geq\ (m-1)\delta_{0}.$
@@
Lemma 2.11 Let $(X,d_{X})$ be a smocked space obtained from $\mathbb{E}^{N}$ by collapsing a collection of compact stitches $\{S_{\alpha}\}$, and let $\pi:\mathbb{E}^{N}\to X$ be the quotient map. Assume that in $\overline{B_{R+1}(0)}$ every stitch has diameter $\leq L$ and distinct stitches are $\delta_{0}$–separated. Then for every $r>0$ there exists $R(r)<\infty$, depending only on $r,L,\delta_{0}$ and $R$, such that
$\pi^{-1}\big(B_{r}^{X}(\pi(0))\big)\subset B_{R(r)}^{\mathbb{E}^{N}}(0).$
In particular, when we assume uniform local bounds on the smocking constants for the sequence $\{X_{k}\}$, this $R(r)$ can be chosen uniformly in $k$.
[…]
@@ 3 Gromov-Hausdorff convergence of smocked spaces
In this section, we will investigate general results on Gromov-Hausdorff convergence, including sufficient and necessary conditions for precompactness and whether or not Hausdorff convergence of the smocking sets implies Gromov-Hausdorff convergence of the smocked spaces. We also prove a general result for tangent cones at infinity. The theory developed in this section thus answers three open questions of Sormani et al.
@@ 3.1 Relation between Hausdorff and Gromov-Hausdorff convergence
An open question for these spaces is whether Hausdorff convergence of the smocking sets $S_{k}$ to $S_{\infty}$ by itself implies Gromov–Hausdorff convergence of the associated smocked spaces $X_{k}$ to $X_{\infty}$ [2]. This is false in general. The key point is that Hausdorff convergence does not by itself control the amount of length that can be collapsed and hidden inside a stitch, which influences distances in the smocked space. This can be demonstrated with an explicit counterexample (the example is one-dimensional, but the construction can easily be generalized).
Example 3.1: Fix the ambient ball radius $R=1$ and the ambient interval $I=[-1,1]\subset\mathbb{R}$. We will construct a sequence of families of disjoint closed intervals $\{S_{\alpha}^{(k)}\}\subset I$ so that the unions $U_{k}:=\bigcup_{\alpha}S_{\alpha}^{(k)}$ converge in the Hausdorff distance to $U:=I$, but the associated smocked spaces $X_{k}$ do not admit a subsequence which converges in the pointed Gromov-Hausdorff sense.
Fix two distinct numbers $L^{\mathrm{odd}},L^{\mathrm{even}}\in(0,2)$. For example, we may use $L^{\mathrm{even}}=1/3$ and $L^{\mathrm{odd}}=2/3$. For each integer $k\geq 1$, set
$N_{k}:=k,\qquad\delta_{k}:=\frac{1}{N_{k}+1}.$
We build $N_{k}$ disjoint small intervals in $I$ so that the maximum length of a gap between adjacent intervals is less than or equal to $\delta_{k}$, hence the Hausdorff distance from $U_{k}$ to $I$ will be $\leq\delta_{k}$ and tends to 0 as $k\to\infty$.
Choose a sequence of total collapsed lengths $L_{k}$ which alternates as follows:
$L_{k}=\begin{cases}L^{\mathrm{even}},&k\ \text{even},\\[4.0pt] L^{\mathrm{odd}},&k\ \text{odd}.\end{cases}$
For each $k$, we distribute the total length $L_{k}$ amongst the $N_{k}$ small intervals approximately equally: set each stitch length $\ell_{j,k}=L_{k}/N_{k}$ for $j=1,\dots,N_{k}$. Since $\ell_{j,k}=L_{k}/N_{k}\to 0$ as $k\to\infty$, the individual stitch diameters tend to 0.
Next, place the $N_{k}$ intervals in $I$ in increasing order so that the gaps between them are all $\leq\delta_{k}$ (this is possible because $N_{k}$ is finite and we may choose the starting position as we wish). Concretely, take the leftmost interval to begin at 1-1 and then place intervals and small gaps alternately so that the number of gaps is $N_{k}+1$ and each gap has length at most $\delta_{k}$. This is possible because
$(N_{k}+1)\delta_{k}=(N_{k}+1)\frac{1}{N_{k}+1}=1,$
so the total length of gaps can be arranged to be less than or equal to 1, leaving enough room for the total stitch length $L_{k}\leq 2$. With minor endpoint adjustments, the construction fits in $I$ for all large $k$. By construction, the maximal gap size in $I\setminus U_{k}$ is $\leq\delta_{k}$, hence $d_{H}(U_{k},I)\leq\delta_{k}\to 0$, so $U_{k}$ converges to $I$ in the Hausdorff metric on compact subsets.
Let $\pi_{k}:\mathbb{R}\to X_{k}$ denote the smocking map which collapse each stitch interval $S_{k}$ to a single point. For any two points, $a<b\in\mathbb{R}$ the smocked space distance $d_{X_{k}}(\pi_{k}(a),\pi_{k}(b))$ is the infimum length of paths joining $a$ to $b$ in $\mathbb{R}$ measured with the convention that any portion lying inside a collapsed stitch contributes zero length. In particular, for the endpoints $1,1\in I$ we have the exact formula
$d_{X_{k}}\big(\pi_{k}(-1),\pi_{k}(1)\big)\;=\;2-L_{k}.$
The interval [1,1][-1,1] has length $2$, and collapsing the union of disjoint subintervals of total length $L_{k}$ reduces the distance by exactly $L_{k}$.
However, since $L_{k}$ alternates between the two distinct values $L^{\mathrm{even}}$ and $L^{\mathrm{odd}}$, the sequence of real numbers given by $d_{X_{k}}$ does not converge (it attains two distinct limit points for the even and the odd subsequences), hence the pointed metric spaces $(X_{k},\pi_{k}(0))$ do not converge to $(X_{\infty},\pi_{\infty}(0))$ in the pointed Gromov–Hausdorff topology.
If we rule out pathological counterexamples, then generally we would expect Hausdorff convergence of the smocking sets to imply Gromov-Hausdorff convergence of the corresponding smocked spaces. A well-behaved example where this is true is provided below.
@@
as required.
Although Hausdorff convergence of the smocking sets does not generally imply Gromov-Hausdorff convergence of the smocking spaces, we can instead prove a natural local replacement: Hausdorff convergence for each bounded ball along with a uniform local bound on the smocking constants is sufficient to imply pointed Gromov-Hausdorff convergence.
Proposition 3.3: Let $\{S_{\alpha}^{(k)}\}_{\alpha\in\mathcal{A}_{k}}$ be the collections of compact stitches in $\mathbb{E}^{N}$ defining smocked spaces $X_{k}=\mathbb{E}^{N}/\!\sim_{k}$ with smocking maps $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ and basepoints $x_{k}=\pi_{k}(0)$. Assume:
  1. (H1)
    (Local Hausdorff convergence) For every $R>0$, the Hausdorff distance between unions of stitches meeting $\overline{B_{R}(0)}$ converges to zero:
    $d_{H}\Big(\bigcup_{\alpha:S_{\alpha}^{(k)}\cap\overline{B_{R}}\neq\varnothing}S_{\alpha}^{(k)},\ \bigcup_{\beta:S_{\beta}^{(\infty)}\cap\overline{B_{R}}\neq\varnothing}S_{\beta}^{(\infty)}\Big)\xrightarrow[k\to\infty]{}0.$
  2. (H2)
    (Local uniform bounds on smocking constants) For every $R>0$, there exist $L_{R}<\infty$, $\delta_{R}>0$ and $K_{R}$ such that for all $k\geq K_{R}$ every stitch intersecting $\overline{B_{R}(0)}$ has diameter $\leq L_{R}$ and distinct such stitches are separated by Euclidean distance at least $\delta_{R}$.
Let $X_{\infty}$ be the smocked space obtained from the limiting stitches $\{S_{\beta}^{(\infty)}\}$, with basepoint $x_{\infty}=\pi_{\infty}(0)$. Then $(X_{k},x_{k})\xrightarrow{\mathrm{pGH}}(X_{\infty},x_{\infty})$ as $k\to\infty$.
Proof: Fix $R>0$ and $\varepsilon>0$. The strategy of the proof is to produce a correspondence between $B_{R}^{X_{k}}(x_{k})$ and $B_{R}^{X_{\infty}}(x_{\infty})$ of distortion $\varepsilon$, for all $k$. By (H1) and (H2) we can choose $\eta>0$ and $K$ so that for $k\geq K$ the finite collections of stitches meeting $\overline{B_{R+1}(0)}$ are in bijection (matched pairwise) and each matched pair is Hausdorff-close by $\eta$. Furthermore, each such stitch has diameter $\leq L:=L_{R+1}$ and distinct stitches are $\geq\delta_{0}:=\delta_{R+1}$ apart. Let $N$ be a uniform bound on the number of stitches meeting $\overline{B_{R+1}(0)}$. This bound exists by (H2).
Define a correspondence $\mathcal{R}\subset X_{k}\times X_{\infty}$ by declaring that $(x_{k},x_{\infty})\in\mathcal{R}$ if and only if one of the following holds:
  1. 1.
    there exists $z\in\overline{B_{R+1}(0)}$ such that $x_{k}=\pi_{k}(z)$ and $x_{\infty}=\pi_{\infty}(z)$; or
  2. 2.
    there is a matched pair of stitches $S_{\alpha}^{(k)}\leftrightarrow S_{\beta}^{(\infty)}$ such that $x_{k}=\pi_{k}(z_{k})$, $x_{\infty}=\pi_{\infty}(z_{\infty})$ for some $z_{k}\in S_{\alpha}^{(k)}$, $z_{\infty}\in S_{\beta}^{(\infty)}$.
The idea of the construction is to pair up points in $X_{k}$ and $X_{\infty}$ that are images of the same preimage $z\in\mathbb{E}^{N}$, and to pair each stitch in $X_{k}$ with its matched limiting stitch in $X_{\infty}$. The projections of $\mathcal{R}$ onto each factor are surjective onto the images of $\overline{B_{R+1}(0)}$ under $\pi_{k}$ and $\pi_{\infty}$, respectively. In particular, $\mathcal{R}$ induces a correspondence between $B_{R}^{X_{k}}(x_{k})$ and $B_{R}^{X_{\infty}}(x_{\infty})$. The margin of +1+1 ensures that every point of the metric balls admits a lift within $\overline{B_{R+1}(0)}$.
Let $(u_{i},v_{i})\in\mathcal{R}$ with lifts $a_{i}\in\overline{B_{R+1}(0)}$ so that $u_{i}=\pi_{k}(a_{i})$ and $v_{i}=\pi_{\infty}(a_{i})$, except for when $a_{i}$ lies in a collapsed stitch, in which case $u_{i}$ (resp. $v_{i}$) is its image under $\pi_{k}$ (resp. $\pi_{\infty}$). Let $\gamma_{k}$ be an $\varepsilon$–minimizing path in $X_{k}$ from $u_{1}$ to $u_{2}$ of length $\ell_{k}\leq d_{X_{k}}(u_{1},u_{2})+\varepsilon$. Lifting $\gamma_{k}$ to $\mathbb{E}^{N}$ produces a concatenation of Euclidean segments joining $a_{1}$ to $a_{2}$. Each time $\gamma_{k}$ crosses a stitch, the lift gains at most $\mathrm{diam}(S_{\alpha}^{(k)})\leq L$ in Euclidean length. Since distinct stitches are $\delta_{0}$–separated, any path of length $\leq 2R+2$ can meet at most $M=M(R,L,\delta_{0})$ of them, hence the lifted path has Euclidean length in $\ell_{k},\ \ell_{k}+ML]$. Performing the same construction in the limit space $X_{\infty}$ gives analogous bounds, and by (H1) the matched stitches are Hausdorff–close, so
$|d_{X_{k}}(u_{1},u_{2})-d_{X_{\infty}}(v_{1},v_{2})|\leq C(R,L,\delta_{0})\,\eta,$
where $\eta$ is the Hausdorff discrepancy of corresponding stitches. It follows that $\mathrm{dis}(\mathcal{R})\to 0$ as $k\to\infty$.
By (H1), each matched pair of stitches is Hausdorff–close by $\eta$ in $\mathbb{E}^{N}$. Combining this with the Euclidean triangle inequality and the previous length estimates gives an additive bound
$\big|d_{X_{k}}(u_{1},u_{2})-d_{X_{\infty}}(v_{1},v_{2})\big|\leq C(R,L,\delta_{0})\,\eta,$
where the constant $C(R,L,\delta_{0})$ depends only on the uniform local geometry and bounds the total contribution from the finitely many stitches that a minimizing path may cross. Choosing $\eta$ small enough so that $C\,\eta<\varepsilon$ yields $\mathrm{dis}(\mathcal{R})<\varepsilon$. Hence, for all sufficiently large $k$,
$d_{GH}\big(B_{R}^{X_{k}}(x_{k}),B_{R}^{X_{\infty}}(x_{\infty})\big)\leq\varepsilon.$
Finally, a standard diagonal argument over $R\to\infty$ gives
$(X_{k},x_{k})\xrightarrow[\;k\to\infty\;]{\mathrm{pGH}}(X_{\infty},x_{\infty}).$
[…]